OpenStudy (anonymous):

An object is launched at 20 m/s from a height of 65 m. The equation for the height (h) in terms of time (t) is given by h(t) = -4.9t2 + 20t + 65. What is the object's maximum height?

OpenStudy (anonymous):

I'm guessing that we want to find the vertex of the function....not exactly sure...what do you think?

OpenStudy (lgbasallote):

it would be when final velocity is zero

OpenStudy (anonymous):

Final velocity is zero when it hits the ground.

OpenStudy (anonymous):

Yeah I believe vertex is what we are trying to find, im not quite sure

OpenStudy (lgbasallote):

what i would do is a little complicated lol...idk how to do this in simple way...but i would take the derivative then equate to 0..solve for t...then sub that value of t to h(t)

OpenStudy (lgbasallote):

i dont think you got to derivative yet..

OpenStudy (anonymous):

That's what I was thinking....

OpenStudy (lgbasallote):

but for the sake of it i'd like to do it :DDD h'(t) = -9.8 t + 20 0 = -9.8t + 20 9.8t = 20 t = 2.04 now sub back in to original equation h(t) = -4.9 (2.04)^2 + 20(2.04) + 65 h(t) = -20.4 + 40.8 + 65 h(t) = 85.4

OpenStudy (lgbasallote):

idk how to do it simply though

OpenStudy (lgbasallote):

finding the maximum using algebra is complicated lol

OpenStudy (anonymous):

thats okay thanks much! :D

OpenStudy (anonymous):

Please don't give answers

OpenStudy (anonymous):

@Romero, if @iceywheniplease7 wanted answers, she'll get it.